The demand and total cost functions of a good are respectively and



Find expressions for TR, (profit) , MR, and MC in terms of Q.

Solve the equation


and hence determine the value of Q which maximizes profit.

Verify that, at the point of maximum profit, MR=MC.


1
Expert's answer
2021-05-26T17:36:25-0400

The demand and total cost functions of a good are


"4P+Q-16=0"

and


"TC=4+2Q-\dfrac{3Q^2}{10}+\dfrac{Q^3}{20}"

respectively.

a)


"4P+Q-16=0"

"P=4-\dfrac{1}{4}Q"


"TR=P\cdot Q=(4-\dfrac{1}{4}Q)\cdot Q=4Q-\dfrac{Q^2}{4}"

"TR=4Q-\dfrac{Q^2}{4}"

"T\pi=TR-TC"

"T\pi=4Q-\dfrac{Q^2}{4}-(4+2Q-\dfrac{3Q^2}{10}+\dfrac{Q^3}{20})"

"T\pi=-4+2Q+\dfrac{Q^2}{20}-\dfrac{Q^3}{20}"


"MR=\dfrac{d(TR)}{dQ}=\dfrac{d}{dQ}(4Q-\dfrac{Q^2}{4})=4-\dfrac{Q}{2}"

"MR=4-\dfrac{Q}{2}"

"MC=\dfrac{d(TC)}{dQ}=\dfrac{d}{dQ}(4+2Q-\dfrac{3Q^2}{10}+\dfrac{Q^3}{20})"

"=2-\dfrac{3Q}{5}+\dfrac{3Q^2}{20}"

b) Solve the equation


"\dfrac{d(T\pi)}{dQ}=0"

"\dfrac{d(T\pi)}{dQ}=\dfrac{d}{dQ}(-4+2Q+\dfrac{Q^2}{20}-\dfrac{Q^3}{20})"

"=2+\dfrac{Q}{10}-\dfrac{3Q^2}{20}"

"2+\dfrac{Q}{10}-\dfrac{3Q^2}{20}=0, Q\geq 0"

"3Q^2-2Q-40=0"

"D=(-2)^2-4(3)(40)=484=22^2"

"Q=\dfrac{2\pm\sqrt{22^2}}{2(3)}=\dfrac{1\pm11}{3}"

"Q_1=-\dfrac{10}{3}, Q_2=4"

Since "Q\geq 0"

If "0\leq Q<4, \dfrac{d(T\pi)}{dQ}>0, T\pi" increases.

If "Q>4, \dfrac{d(T\pi)}{dQ}<0, T\pi" decreases.

The function "T\pi" has a local maximum at "Q=4."

Since the function "T\pi" has the only extremum for "Q\geq 0," then the function "T\pi" has the absolute maximum at "Q=4" for "Q\geq 0."


c) "Q=4"


"MR(4)=4-\dfrac{4}{2}=2"

"MC(4)=2-\dfrac{3(4)}{5}+\dfrac{3(4)^2}{20}=2"

Then at the point of the maximum profit "(Q=4)"


"MR(4)=2=MC(2)"